# LayerPano3D: Layered 3D Panorama for Hyper-Immersive Scene Generation

### v1

@@ 4.4 Quantitative Comparison
3D Panoramic Scene Reconstruction. Following [2], we adopt two widely-used non-reference image quality assessment metrics, NIQE [61] and BRISQUE [62], to evaluate the capability of free trajectory rendering, i.e. the novel view quality along an immersive navigation path in the 3D scene. Furthermore, we also measure the 3D reconstruction quality on training views via classic reconstruction metrics: PSNR, SSIM and LPIPS [63]. In addition, we render four orthogonal views, each distributed at a fixed FOV ($90^{\circ}$), elevation ($0^{\circ}$) and four equidistant azimuths ($0^{\circ},90^{\circ},180^{\circ},270^{\circ}$) respectively, to predict perspective field. Building upon  [64], we independently predict camera parameters and evaluate the horizon tilt problem by calculating the Roll-Mean and Roll-Var, i.e. the mean and variance of the elevation angles across these views. As shown in Table II, our method surpasses the existing methods in both novel view quality metrics (NIQE and BRISQUE) and 3D reconstruction metrics (PSNR, SSIM, and LPIPS). Furthermore, we conducted another user study to further evaluate the quality of the generated 3D panoramic scene from two aspects: 1) $360^{\circ}\times 180^{\circ}$ view consistency and 2) free path rendering quality. For the first aspect, we render 60 frames in a 360-degree view at the 0-degree and 45-degree elevation respectively for evaluation. For the second aspect, we select the same trajectory as in Figure 7 to render navigation videos for evaluation. We invite 43 users including graduate students that expertise in 3D and average users to rank the 40 results from 4 methods individually. The average ranking is shown in Table II. Our LayerPano3D achieves the best performance in both $360^{\circ}\times 180^{\circ}$ view consistency and free path rendering quality among all four approaches.
Figure 9: Ablation on the Circular Blending Strategy. The circular blending strategy ensures the continuity between the leftmost and rightmost sides of the panorama, which preserves the crucial characteristic of 360-degree panoramas.
Figure 10: Ablation on the Gaussian Selector. Our Gaussian Selector resolves the optimization conflicts between layers by re-activating the Gaussians from the previous layer that blocks new assets. With the Gaussian Selector, the merged Gaussians are optimized to faithfully reconstruct the ground-truth panorama views.
4.5 Analysis and Ablative Study
Ablation on Circular Blending. Figure 9 demonstrates that the circular blending strategy effectively maintains geometric continuity, contributing to the generation of a seamless 360-degree panorama. Without this design, the rightmost and leftmost sides of the panorama could form inconsistent patterns, resulting in a discontinuous panorama.
Analysis on Layer Completion Inpainting. We discuss the effectiveness of our fine-tuned panorama inpainter in layer completion. We compare the inpainting results among three approaches: LaMa [42], Stable Diffusion inpainting model and our proposed inpainter. As illustrated in Figure 8, LaMa produces inconsistent texture and blurry artifacts at large-scale inpainting. Stable Diffusion tends to produce distorted new elements due to the domain gap between perspective and panoramic images. In contrast, thanks to the panorama finetuning, our module delivers clean inpainting results with coherent and plausible structures in the masked regions.
@@ 4.5 Analysis and Ablative Study
Ablation on Gaussian Selector. Our Gaussian selector is proposed to select the part of Gaussians that appears in the front of newly added scene assets. By selecting these Gaussians and re-activating them in the optimization, the model achieves accurate appearance and geometry at the current layer. As shown in Figure 10, the leftmost column is the scene Gaussians at layer 0. When adding the building assets at the first layer, the sky Gaussians from the previous layer partially block the building assets (right column). After using the Gaussian selector to select and optimize the sky Gaussians, these Gaussians learn to either be translucent and pruned for low opacity or move to be a part of the building assets. Therefore in the middle column, we observe a consistent scene with no obvious blockage of the new building assets thanks to the Gaussian Selector.
@@
Figure 11: Analysis on Panorama Rendering at Off-center Viewpoints. We show that LayerPano3D ’s robustness in rendering $360^{\circ}\times 180^{\circ}$ consistent panorama at various viewpoints within a large range of space in the generated scene. Compared with our single-layer panorama baseline, LayerPano3D renderings exhibit high-quality content without any holes or gaps from occlusion.
@@
Analysis on Panorama Renderings at Off-center Viewpoints. In Figure 11, we demonstrate that LayerPano3D is robust to render consistent panorama images at various locations besides the original camera location in the center. We sample four camera locations on circular trajectories on the hemisphere centered at the origin and render 24 views at ($-45^{\circ},0^{\circ},45^{\circ}$) elevation to compose new panorama images. By evaluating panorama renderings at new viewpoints, we show that our generated panoramic scene is $360^{\circ}\times 180^{\circ}$ consistent and enclosed, robust to various viewpoints at any angle. Compared to the single-layered 3D panorama, our multi-layered 3D panorama exhibits no gaps or holes from the scene occlusion, demonstrating our capability for free 3D exploration in the generated scenes.
5 Conclusion
@@
In this paper, we propose LayerPano3D, a novel framework that generates hyper-immersive panoramic scene from a single text prompt. Our key contributions are two-fold. First, we propose the text-guided anchor view synthesis pipeline to generate detailed and consistent reference panorama. Second, we pioneer the Layered 3D Panorama representation to show complex scene hierarchies at multiple depth layers, and lift it to Gaussians to enable free 3D exploration. Extensive experiments show the effectiveness of LayerPano3D in generating $360^{\circ}\times 180^{\circ}$ consistent panorama at various viewpoints and enabling immersive roaming in 3D space. We believe that LayerPano3D holds promise to advance high-quality, explorable 3D scene creation in both academia and industry.
@@
Figure 12: Additional results of  LayerPano3D on Diverse Generation. LayerPano3D generates various hyper-immersive Layered 3D Panorama that cover cityscape, landscape and indoor environments.

### v2

@@ 4.4. Quantitative Comparison
Following (Zhou et al., 2024b), we adopt non-reference image quality assessment metrics, NIQE (Mittal et al., 2012b) and BRISQUE (Mittal et al., 2012a), to evaluate novel view quality along scene navigation paths. We also follow (Zhou et al., 2024b) to measure the rendering quality with PSNR, SSIM and LPIPS (Zhang et al., 2018). In terms of geometry evaluation, we render 4 orthogonal views ($90^{\circ}$ FOV, $0^{\circ}$ elevation and { $0^{\circ},90^{\circ},180^{\circ},270^{\circ}$ azimuths) and predict the Pitch-Mean and Pitch-Var (mean and variance of the elevation angles) with (Veicht et al., 2024) to evaluate whether the scenes are upright. As shown in Tab. 2, our method surpasses the existing methods in both novel view quality metrics (NIQE and BRISQUE), 3D reconstruction metrics (PSNR, SSIM, LPIPS) and while ensuring the upright panoramic scene. Furthermore, we conduct another user study for 3D panoramic scene evaluation from two aspects: 1) $360^{\circ}\times 180^{\circ}$ view consistency and 2) novel path rendering quality. For the first aspect, we render 60 frames to cover 360-degree view at the 0-degree and 45-degree elevation respectively for evaluation. For the second aspect, we use the same trajectory as in Fig. 5 to render navigation videos for evaluation. We invite 52 users including graduate students that expertise in 3D and average users to rank the 40 results from 4 methods. The average ranking is shown in Tab. 2. Our LayerPano3D achieves the best performance in both $360^{\circ}\times 180^{\circ}$ view consistency and novel path rendering quality among all four approaches.
Figure 6. Qualitative comparisons in Panorama Generation. LayerPano3D demonstrates superior capability in generating high-quality outputs with precise alignment to text prompt, outperforming other methods in fidelity and input adherence.
4.5. Analysis and Ablative Study
In this section and supp., we show the analysis and ablation on the Gaussian Selector (Sec. 4.5.1), Multi-layer design (single vs. multi; Sec. 4.5.2), Layer Gaussians representation (supp.), layer inpainting (supp.) and 3DGS optimization efficiency (supp.).
4.5.1. Ablation on Gaussian Selector.
@@ 4.5. Analysis and Ablative Study
Our Gaussian selector is proposed to select the part of Gaussians that appears in the front of newly added scene assets. By selecting these Gaussians and re-activating them in the optimization, the model achieves accurate appearance and geometry at the current layer. As shown in Fig. 7, the leftmost column is the scene Gaussians at layer 0. When adding the building assets at the first layer, the sky Gaussians from the previous layer partially block the building assets (right column). After using the Gaussian selector to select and optimize the sky Gaussians, these Gaussians learn to either be translucent and pruned for low opacity or move to be a part of the building assets. Therefore in the middle column, we observe a consistent scene with no obvious blockage of the new building assets thanks to the Gaussian Selector.
Figure 7. Ablation on the Gaussian Selector. With the Gaussian Selector, the merged Gaussians are optimized to faithfully reconstruct the ground-truth panorama views.
@@
Figure 8. Analysis on Panorama Rendering at Off-center Viewpoints. Compared with the single-layer variant, LayerPano3D render $360^{\circ}\times 180^{\circ}$ consistent panorama at various off-center viewpoints without any holes or gaps from occlusion.
4.5.2. Analysis on Panorama Renderings at Off-center Viewpoints.
@@
In Fig. 8, we demonstrate that LayerPano3D is robust to render consistent panorama images at various locations besides the original camera location in the center. We sample four camera locations on circular trajectories on the hemisphere centered at the origin and render 24 views at ($-45^{\circ},0^{\circ},45^{\circ}$) elevation to compose new panorama images. By evaluating panorama renderings at new viewpoints, we show that our generated panoramic scene is $360^{\circ}\times 180^{\circ}$ consistent and enclosed, robust to various viewpoints at any angle. Compared to the single-layered 3D panorama, our multi-layered 3D panorama exhibits no gaps or holes from the scene occlusion, demonstrating our capability for larger-range, complex 3D exploration in the generated scenes.
5. Conclusion
@@
In this paper, we propose LayerPano3D, a novel framework that generates hyper-immersive panoramic scene from a single text prompt. Our key contributions are two-fold. First, we propose the text-guided anchor view synthesis pipeline to generate detailed and consistent reference panorama. Second, we pioneer the Layered 3D Panorama representation to show complex scene hierarchies at multiple depth layers, and lift it to Gaussians to enable large-range 3D exploration. Extensive experiments show the effectiveness of LayerPano3D in generating $360^{\circ}\times 180^{\circ}$ consistent panorama at various viewpoints and enabling immersive roaming in 3D space. We believe that LayerPano3D holds promise to advance high-quality, explorable 3D scene creation in both academia and industry.